A batter strikes a baseball (mass = 145 g) which is moving at 40 m/s such that it leaves the bat at a speed of 80 m/s. If the bat is in contact with the ball for 1.10 ms, what is the magnitude of the peak force on the ball?
Answer #1

Impulse = m* ΔV
Impulse = F*Δt

So We can say ,
F = (m*ΔV) / (Δt)

ΔV = 40 - (-80) = 120 m/s
Δt = 1.10 * 10^-3 s
m = 145 * 10^-3 Kg

F = (m*ΔV) / (Δt)
F = (145*10^-3 * 120)/ (1.1 * 10^-3) N
F = 1.58 * 10^4 N
Magnitude of the peak force on the ball, F = 1.58 * 10^4 N